OpenStudy (anonymous):

May has 10 fewer marbles than Albert May has 5 more marbles than Jenny Albert gives 4 marbles to Jenny May gives antoher 6 marbles to Jenny Jenny has 13 marbles now how much may and ablert have at first

OpenStudy (konradzuse):

May = x -10 Albert = x Jenny = x - 15 Albert = x - 4 May = x - 16 Jenny = (x - 15 + 6 + 4) = 13

OpenStudy (anonymous):

what one did they start out with?

OpenStudy (konradzuse):

doesn't matter really, we only needed the original info to determine an X.

OpenStudy (anonymous):

so it dosent matter which one they started out with?

OpenStudy (konradzuse):

it does, but we need to find out x from the end point... Do you see it?

OpenStudy (anonymous):

no can you explain

OpenStudy (konradzuse):

Jenny = (x - 15 + 6 + 4) = 13

OpenStudy (konradzuse):

we know Jenny = 13, so we can figure out x from this.

OpenStudy (anonymous):

oh okay i get it

OpenStudy (anonymous):

thank you!

OpenStudy (konradzuse):

show me what you did.